Telescope and AGw4 Commissioning at LDG - 12-13 May 2009 MST

Observers:: J. Hill (mtn) , A. Rakich (Tucson)
Telescope Operator: A. Ceranski
Software Support: C. Biddick (phone)
Telescope Support: T. Shih, J. Little

Summary

We applied Jesper's new AGW transform (the first from measured AGW4 data on Sunday night), and it seemed to work OK.

We found a backdoor to a bug in GCS we thought we had fixed last week, so Chris made another patch. After this initial fight with GCS, the telescope was extremely cooperative.

We took more AGW transformation data with the new transformation on BS9152 and StoneM. Conditions were acceptable, but not excellent. We also took cubes to measure the rotator center on IRTC.

We took offset data with the new transformation.

We took some photometric data on M92 to replace the especially ratty ones from last week.

This ends the May 2009 campaign for Direct Gregorian Commissioning.

Preparations

Gene was working on the LDG rotator this morning. Chris sent a patch to PMC related to the integrator on the temperature control. Butch and friends were working on the labyrinth seal for the rotating building.

Jesper sent a refined AGW transformation based on Sunday's data. "A simple orthogonal transform gives an RMS of 42 and 57 microns in X and Y respectively. If I include the pinhole offset I get RMS values of 27microns in both X and Y, so this looks excellent. The pinhole offset is fitted at 0.86mm which I believe is very reasonable, so this should be a physically acceptable solution. I have adopted the mean (x,y) value on the IRTC as the reference point in lack of a true rotation axis position there, but that should only be a matter of an offset.

I have left a modified configuration file with the name oacontrol.conf.agw4trans on agw-control which you can install when you want.

I have changed the pcsmx flag from -1 to +1 and the m_xgp flag from +1 to -1 to be consistent with the fit. I hope that I have not gotten tangled in the signs here..."

18:45 Aaron had trouble getting LDG rotator running. It worked after we stopped and restarted MCSPU. Apparently you have to do this whenever anyone (Gene) has disconnected/reconnected the fiber.

19:00 John crashed TCS with a readWFSCam command. It tried to selectaGW AIP_L and then crashed when AGW unit 0 didn't exist. Called Chris Biddick. The reason this didn't crash during the night last night was that it only selects a new AGW when NOT guiding, so IDL was able to use it OK.

Details

Open

19:12 Open Dome. T=9.8 degC D=-14.2 degC Wind = 7 m/s, Sky photometric.

19:20 Set Ki for right mirror ventilation to +0.01.

19:29 pointto M5_1246 Tel="ACQUIRE" PAM- gs=0 - hits the star within 6 arcsec.

19:34 pointto M5_1246 Tel="ACQUIRE" PAM- gs=0 - hits the same star 3 arcsec different in IE (down on the image)

19:35 pointto M5_1246 Tel="ACQUIRE" PAM- gs=0 - hits the same star 2 arcsec different in IE (down on the image)

Ztemp changed from 1.228 to 1.231 mm (implying a similar change in y) between these two.

19:37 pointto M5_1246 Tel="ACQUIRE" PAM- gs=0 - hits the same place on fourth try -Adjust IE=-38 CA=+13

19:45 pointto BS9145 Tel="ACQUIRE" PAM- gs=0

Chris has patched readWFSCam and readGuideCam to not ever call selectAGW AIP_L.cfg. You must still issue the appropriate selectAGW command when you start GCS.

stopAGW -u 1, startAGW -u 1, selectAGW............. (This probably was not necessary as the problem turned out to be doing a short exposure while the dome was light.)

Exposures shorter than 400 ms didn't work. readGuideCam -e 400, but worked for readWFSCam down to 1 ms. See Issue #2175.

20:06 pointto BS9145 Tel="ACQUIRE" PAM- gs=0 - probe hadn't been homed since we restarted.

20:09 pointto BS9145 Tel="ACQUIRE" PAM- gs=0 CA=5

20:15 pointto BS9145 Tel="ACTIVE" PAM+ gs=8 -- seeing is 0.5 arcsec!

20:20 copy oacontrol.conf.agw4trans to replace oacontrol.conf.20090506b as /etc/oacontrol/oacontrol.conf

20:23 stopGuiding; stop and start oacserver. (sudo /etc/init.d/oacsrv start, selectAGW.........).

20:25 pointto BS9145 Tel="ACTIVE" PAM+ gs=8 PA=INDEF=344.29

20:30 pointto BS9145 Tel="ACTIVE" PAM+ PA=0 gs=9

Focus Test

20:33 stopWFS

focus=31.0 , and probe is at 25.872 off-axis.

20:35 cl < home$focus_seq.cl
irtc.20090513.033549.fits cube 0101 +0.045 mm
irtc.20090513.033734.fits cube 0102 +0.030 mm
irtc.20090513.033912.fits cube 0103 +0.015 mm
irtc.20090513.034048.fits cube 0104 0.0
irtc.20090513.034228.fits cube 0105 -0.015 mm
irtc.20090513.034406.fits cube 0106 -0.030 mm
irtc.20090513.034546.fits cube 0107 -0.045 mm
(no sky for this one)

By eye, focus is near the second image (or between 2 and 3) so say dZ=+0.025 mm.

   irc Move 2. 0. 0. 0. 0. 0. 1 ABS M2 0 left

...Done! (success)

ERROR: package `1' not found
     called as: `cl ()'
There was a stray decimal point in an unfortunate location in my script. Clear instrument offsets in PSF to clean up after the crashed script.

20:48 resend active preset

20:55 Focus to 33.4 on-axis in AIP_IRTC_DG_L.cfg, selectAGW.........

20:56 resend active preset, now off-axis focus is 28.764

Andrew says the axial magnifcation is 169. 1 mm Z-motion of M2 move the focal plane by 169 mm.

IIFGUI is confused that WFS is Paused even though GCS is sending corrections. There must have been an interrupted offset somewhere.

23:01 cl < home$focus_seq.cl
irtc.20090513.040233.fits cube 0108 +0.045
irtc.20090513.040421.fits cube 0109 +0.030
irtc.20090513.040604.fits cube 0110 +0.015
irtc.20090513.040744.fits cube 0111 0.0
irtc.20090513.040926.fits cube 0112 -0.015
irtc.20090513.041106.fits cube 0113 -0.030
irtc.20090513.041247.fits cube 0114 -0.045
irtc.20090513.041416.fits cube 0115 sky

Need to analyze this set of cubes for a precise answer.

AGW Transformation Data

21:17 pointto BS9152 Tel="ACTIVE" PAM+ PA=0 gs=0

21:23 Setting none mode from mirror panic. (It would help if the message said why it went to none mode.) Lost 10 minutes.

Reduce Ki to 0.003 on DX Mirror Ventilation.

21:29 pointto BS9152 Tel="ACTIVE" PAM+ PA=0 gs=0

21:36 cl < home$tran_BS9152_00.cl - seeing is 0.55 to 0.75 in this dataset. Arbitrator is in 'split' mode.
transtar (2) irtc.20090513.043838.fits cube 0116
transtar (4) irtc.20090513.044100.fits cube 0117
transtar (7) irtc.20090513.044314.fits cube 0118
transtar (10) irtc.20090513.044543.fits cube 0119
transtar (14) irtc.20090513.044806.fits cube 0120
transtar (15) irtc.20090513.045018.fits cube 0121
transtar (18) irtc.20090513.045237.fits cube 0122
transtar (19) couldn't move to gs19, so image is not useful, except as the sky cube
transtar (2) irtc.20090513.045647.fits cube 0124

Wind is up to 12 m/sec.

21:59 cl < home$tran_BS9152_90.cl
transtar (12) irtc.20090513.050023.fits cube 0125
transtar (10) irtc.20090513.050240.fits cube 0126
transtar (14) irtc.20090513.050443.fits cube 0127
transtar (15) irtc.20090513.050647.fits cube 0128
transtar (18) irtc.20090513.050853.fits cube 0129
transtar (19) irtc.20090513.051056.fits cube 0130
transtar (11) irtc.20090513.051306.fits cube 0131
transtar (13) irtc.20090513.051510.fits cube 0132
transtar (16) irtc.20090513.051716.fits cube 0133
transtar (17) irtc.20090513.051921.fits cube 0134
transtar (12) irtc.20090513.052129.fits cube 0135

22:24 cl < home$tran_BS9152_180.cl seeing 0.5 to 1 in this data - highly variable
transtar (12) irtc.20090513.052557.fits cube 0136
transtar (11) irtc.20090513.052820.fits cube 0137
transtar (13) irtc.20090513.053030.fits cube 0138
transtar (16) irtc.20090513.053239.fits cube 0139
transtar (17) irtc.20090513.053450.fits cube 0140
transtar (9) irtc.20090513.053638.fits cube 0141
transtar (5) irtc.20090513.053826.fits cube 0142
transtar (1) irtc.20090513.054015.fits cube 0143
transtar (12) irtc.20090513.054230.fits cube 0144

Offset Tests

22:44 pointto BS9152 Tel="ACTIVE" PAM+ PA=0 gs=2 - star not found

22:48 pointto BS9152 Tel="ACQUIRE" PAM+ PA=INDEF gs=1 - Adjust IE=-55 CA=+8

22:51 pointto BS9152 Tel="ACTIVE" PAM+ PA=0 gs=2

22:54 dice5 off1=6.5 off2=5.0 coord="DETXY" exptime=1.2 Nexpo=30 loopN=10
starting with irtc.20090513.055515.fits cube 0145. Seeing was 0.45 - 0.85 at the start.

This dice5 sequence had unfortunate timing. It has just enough time to take WFS image, but not enough time to apply it to PSF. So the result is that active optics is only being applied infrequently.

23:51 pointto BS9152 Tel="ACTIVE" PAM+ PA=330 gs=2

23:54 dice5 off1=6.0 off2=5.0 coord="RADEC" exptime=1.2 Nexpo=40 loopN=10
starting with irtc.20090513.065448.fits cube 0195. Seeing was 0.65 - 0.8 at the start.

Rotator Center measurement on IRTC

01:02 takeseq 1.2 100 while rotating rotator 0-360. irtc.20090513.080309.fits cube 0245
Cube contains points before and after rotation to check the rotator scale. Telescope is tracking open-loop.

01:05 takeseq 1.2 100 back the other way. irtc.20090513.080546.fits cube 0246

M92 Photometric Images

01:08 pointto M92astrom Tel="ACTIVE" PAM+ PA=0 gs=0 17:16:47.9000 43:06:21.900 RADEC 2000.0
I didn't try to verify that this was the right star, but it probably isn't more than 10 arcsec off based on the previous and following observations.

Images in /Repository/AGW_Data/090513w/

stopGuiding

"r"=2 guider0007-0011 - five 10 sec exposures

"v"=1 guider0012-0016 - five 10 sec exposures

01:19 pointto M92astrom Tel="ACTIVE" PAM+ PA=0 gs=0

"b"=0 guider0017-0021 - five 20 sec exposures

01:27 pointto M92astrom Tel="ACTIVE" PAM+ PA=0 gs=0

stopGuiding

offset 0 8

b=0 guider0022-0024 three 20 sec

v=1 guider0025-0027 three 10 sec

r=2 guider0028-0033 six 10 sec

More AGW Transformation Data

01:36 pointto StoneMagw Tel="ACTIVE" PAM+ PA=0 gs=0 - seeing about 0.9 arcsec

01:43 pointto StoneMagw Tel="ACTIVE" PAM+ PA=0 gs=51

01:45 cl < home$tran_StoneM_00.cl - abort because sequence is too short to allow WFS These were 0.5 x 60
epar transtar to set the sequence parameters.

transtar (39)  irtc.20090513.084703.fits cube 0247
transtar (40) irtc.20090513.084910.fits cube 0248

01:50 cl < home$tran_StoneM_00.cl PA=0 These are 0.5 x 80 - seeing varies 0.8 to 1.3 arcsec
Note that successive acquisition images should give a good estimate of open-loop tracking drift. By eyeball it is about 6 arcsec per hour.

transtar (39) irtc.20090513.085117.fits cube 0249
transtar (40) irtc.20090513.085323.fits cube 0250
transtar (51) irtc.20090513.085540.fits cube 0251 
transtar (72) irtc.20090513.085748.fits cube 0252
transtar (68) irtc.20090513.085956.fits cube 0253
transtar (58) irtc.20090513.090200.fits cube 0254
transtar (52) irtc.20090513.090408.fits cube 0255
transtar (48) irtc.20090513.090619.fits cube 0256
transtar (44) irtc.20090513.090824.fits cube 0257
transtar (43) irtc.20090513.091034.fits cube 0258
transtar (38) irtc.20090513.091245.fits cube 0259
transtar (37) irtc.20090513.091456.fits cube 0260
transtar (35) irtc.20090513.091705.fits cube 0261 - something wrong here
transtar (26) irtc.20090513.091911.fits cube 0262
transtar (20) irtc.20090513.092114.fits cube 0263
transtar (13) irtc.20090513.092322.fits cube 0264
transtar (18) irtc.20090513.092535.fits cube 0265
transtar (51) irtc.20090513.092747.fits cube 0266

02:29 takeseq 0.5 80 - irtc.20090513.092854.fits cube 0267 - forgot to offset so this repeats cube 0266

02:30 offset 30 30 to get blank sky

02:31 takeseq 0.5 80 - irtc.20090513.093033.fits cube 0268

02:32 cl < home$tran_StoneM_90.cl PA=90 These are 0.5 x 80. Seeing is 0.8-1 arcsec.

transtar (55) IRTC image failed, but has a getdat entry.
transtar (53)  irtc.20090513.093424.fits cube 0269
transtar (62) irtc.20090513.093633.fits cube 0270
transtar (64) irtc.20090513.093842.fits cube 0271
transtar (69) irtc.20090513.094052.fits cube 0272
transtar (51) irtc.20090513.094307.fits cube 0273
transtar (58) irtc.20090513.094519.fits cube 0274
transtar (68) irtc.20090513.094727.fits cube 0275 - crossing the meridian
transtar (72) irtc.20090513.094941.fits cube 0276
transtar (85) irtc.20090513.095151.fits cube 0277
transtar (83) irtc.20090513.095402.fits cube 0278
transtar (76) irtc.20090513.095616.fits cube 0279
transtar (79) irtc.20090513.095824.fits cube 0280
transtar (67) irtc.20090513.100039.fits cube 0281 
transtar (78) irtc.20090513.100250.fits cube 0282
transtar (74) irtc.20090513.100456.fits cube 0283
transtar (66) irtc.20090513.100702.fits cube 0284
transtar (60) irtc.20090513.100910.fits cube 0285
transtar (63) irtc.20090513.101118.fits cube 0286
transtar (61) irtc.20090513.101326.fits cube 0287
transtar (54) irtc.20090513.101533.fits cube 0288
transtar (49) irtc.20090513.101741.fits cube 0289
transtar (59) irtc.20090513.101955.fits cube 0290
transtar (50) irtc.20090513.102207.fits cube 0291
transtar (55) irtc.20090513.102412.fits cube 0292

03:02 Increase Kp from 1.25 to 1.5 on both mirror cells. Both have the normal offset of -0.9. Only DX has the integrator of 0.003. In a night of rapidly falling temperature (-0.5 deg/hour), SX was lagging 0.4 degrees behind, while DX was a bit overcooled with the integrator causing some oscillation.

03:20 Set both SX and DX mirror ventilation to Offset =-0.9 Kp=1.6 Ki=0.002

03:25 cl < home$tran_StoneM_180.cl PA=180 These are 0.5 x 80. Seeing is 0.7 - 1 arcsec. Guide exposures are 2.0 sec. Effective guide rate is 0.2 Hz, but cadence may be influenced by the frequent presetting.

transtar (55) irtc.20090513.102638.fits cube 0293
transtar (53) irtc.20090513.102844.fits cube 0294
transtar (62) irtc.20090513.103050.fits cube 0295
transtar (76) irtc.20090513.103256.fits cube 0296
transtar (79) irtc.20090513.103512.fits cube 0297
transtar (67) irtc.20090513.103717.fits cube 0298
transtar (78) irtc.20090513.103926.fits cube 0299
transtar (74) irtc.20090513.104133.fits cube 0300
transtar (66) irtc.20090513.104339.fits cube 0301
transtar (60) irtc.20090513.104548.fits cube 0302
transtar (63) irtc.20090513.104755.fits cube 0303
transtar (61) irtc.20090513.105002.fits cube 0304
transtar (54) irtc.20090513.105206.fits cube 0305
transtar (49) irtc.20090513.105411.fits cube 0306
transtar (59) irtc.20090513.105620.fits cube 0307
transtar (50) irtc.20090513.105820.fits cube 0308
transtar (45) irtc.20090513.110028.fits cube 0309
transtar (36) irtc.20090513.110237.fits cube 0310 - has a faint companion
transtar (46) irtc.20090513.110500.fits cube 0311
transtar (41) irtc.20090513.110702.fits cube 0312
transtar (34) irtc.20090513.110911.fits cube 0313
transtar (29) irtc.20090513.111118.fits cube 0314 - has several faint companions = bad coords
transtar (32) irtc.20090513.111329.fits cube 0315
transtar (30) irtc.20090513.111538.fits cube 0316
transtar (25) irtc.20090513.111741.fits cube 0317
transtar (31) irtc.20090513.111953.fits cube 0318
transtar (27) irtc.20090513.112201.fits cube 0319
transtar (21) irtc.20090513.112414.fits cube 0320
transtar (16) irtc.20090513.112622.fits cube 0321
transtar (11) irtc.20090513.112833.fits cube 0322
transtar (9) irtc.20090513.113000.fits cube 0323 - couldn't move probe, telescope open loop - don't use
transtar (5) irtc.20090513.113117.fits cube 0324 - couldn't move probe, telescope open-loop - don't use
transtar (55) irtc.20090513.113325.fits cube 0325

04:01 setGuideRate to 1 Hz, but that doesn't make it go faster. There is some tricky interaction between guide rate and exposure time and overheads.

04:35 cl < home$tran_StoneM_00.cl PA=0 Seeing is 0.6 to 0.8 arcsec.
transtar (39) irtc.20090513.113626.fits cube 0326
transtar (40) irtc.20090513.113838.fits cube 0327
transtar (51) irtc.20090513.114048.fits cube 0328
transtar (72) irtc.20090513.114302.fits cube 0329
transtar (68) irtc.20090513.114513.fits cube 0330
transtar (58) irtc.20090513.114719.fits cube 0331
transtar (52) irtc.20090513.114931.fits cube 0332
transtar (48) irtc.20090513.115137.fits cube 0333
transtar (44) irtc.20090513.115344.fits cube 0334
transtar (43) irtc.20090513.115556.fitscube 0335 - unguided as twilight is upon us

IRTC Data in /Repository/IRTC_Data/090513/

04:52 Guide star is landing within 2 arcsec of hotspot, and we haven't adjusted IE/CA in 6 hours, although it did drift out to 6 arcsec 2 hours ago.

Close

05:00 Close Dome for daylight. T=6.3 degC D=-12 degC, Wind 4 m/sec. Photometric.

Stopped IRTC program. stopAGW -u 1.

AGW4 Camera Saturation Levels

15:20 Verified that both the Guider and the WFS camera in AGW4 have saturation levels that go to 65K counts. This isn't any special photometric statement, but simply empirical confirmation that the readout is that deep (in contrast to the AGW3 WFS camera which saturates at 30k counts).

Task List for Tonight (May 13 UT) and Beyond

LDG specific

  • Measure AGW4-IRTC2 focus (done May 13 UT, pending data reduction)
  • More Transformation Data to test refined transformation (done May 13 UT, pending data reduction)
  • Offset tests with refined transformation (some done May 13 UT)
  • Check DG rotator scale
  • Collect better photometric data on M92 (done 13 May UT)
  • EF Data (intra/extrafocal pupils) to confirm Active Optics performance

Bad weather

  • Measure WFS and Guider camera saturation levels for AGW4 (done May 13 UT)
  • Work on the transistion between AGW3 and AGW4 - it has had problems with oacserver and GCS
  • Calibrate timing for IRTC sequences. Issue 2173 (irc wait time versus arrival time on disk)
  • Enhance transtar.cl to print a log file with IRAF commands to centroid the cubes.

General (could be either LDG or LFBG)

  • Guiding tests versus magnitude
  • Guiding scale factor (waiting on analysis of the previous data from last week)
  • Investigate the adjustment of exposure times in GCS

You can never have too much of ......(could be either LDG or LFBG)

  • Vibration Data on Bright Stars
  • Open-Loop Tracking Data
  • Closed-Loop Tracking Data

-- JohnHill - 12 May 2009
Topic attachments
I Attachment Action Size Date Who Comment
AIP_IRTC_DG_L.cfgcfg AIP_IRTC_DG_L.cfg manage 8 K 13 May 2009 - 22:43 JohnHill GCS configuration file for IRTC at LDG
focus_seq.clcl focus_seq.cl manage 3 K 13 May 2009 - 04:12 JohnHill  
Topic revision: r6 - 13 May 2009, JohnHill
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback